• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Olivier/Brad/Peter

    Log entry time 11:25:38 on March15,2005

    Entry number 141579

    keyword=T7 trigger rate decreased

    We went down the hall to look at T7 trigger, with the idea of increasing the threshold of this trigger to lower the rate. Investigating this, we found some interesting stuff.

    1) What I THOUGHT T7 was: the analog signal from the left PMTs are fed into a summing module (summed bars are on top of each other, not following paths from the target), in groups of 8. All the summed outputs are fed into a discriminator, with a threshold of 140 mV. This discriminated outputs are then put through a majority logic unit, effectively acting as an OR of the summed signals. To some extent, this is equivalent to an OR of all 120 left PMT, with a 140 mV threshold on the input (amplified 10 times) signal.

    2) What T7 actually IS: the same as above, but added in the input of the majority logic unit was a OR of the summed inputs of 4 of the "refreshing" discrimators (Lecroy 4413) that are feeding the TDCs. That was effectively acting as an OR of 64 channels with an 50 mV threshold, thus including a lot of background.

    What we did: we looked at cosmic rates for different setups:

    - setup 2) (old setup): 6 kHz
    - then we removed the extra 64 channels, and looked at the rate: 3.2 kHz
    - then we increased the threshold on the summed signals from 140 to 300 mV -> rate went down to 300 Hz.

    So this is the new setup. And the T7 rate with beam on should be a lot more "workable"!